# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

### v1

@@ B.3. Returning to the original variable
(B.12) $\begin{split}J(w)+&\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}|rw|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{2}_{1}\varphi|rg|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}|rD_{i}(g)|^{2}\,dt\\ &\leq C^{\prime}_{s_{2},\epsilon_{2}}\left(\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}|z|^{2}\,dt+\mathbb{E}\int_{Q}(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi)|\nabla_{\gamma}\psi|^{2}|dz|^{2}\right.\\ &\left.+\sum_{i=1}^{n}\left[\mathbb{E}\int_{Q}s\lambda^{2}\varphi|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}|D_{i}(dz)|^{2}\right]\right)\end{split}$
where $J(w)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}|rD_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{2}\varphi|rA_{i}D_{i}w|^{2}\,dt$ and $w$ is solution of $dw+\sum_{i=1}^{n}D_{i}(\gamma_{i}w)dt=fdt+gdB(t)$ with $w=0$ on $\partial Q$.
@@
Recalling $w=\rho z$, we have $D_{i}w=A_{i}\rho\,D_{i}z+D_{i}\rho\,A_{i}z$. Therefore, by Lemma A.1 and A.3, for each $i\in\llbracket 1,n\rrbracket$ it holds that
(B.13) $\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}|rD_{i}w|^{2}\,dt\leq\;&\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\\ &\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\varphi^{3}\lambda^{4}|A_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|A_{i}z|^{2}\,dt.\end{split}$
@@
(B.14) $\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\varphi^{3}\lambda^{4}|A_{i}z|^{2}\,dt\leq\mathbb{E}\int_{Q}s^{3}\varphi^{3}\lambda^{4}|z|^{2}\,dt+\mathbb{E}\int_{Q}s\mathcal{O}_{\lambda}((sh)^{2})|z|^{2}\,dt.$
@@
(B.15) $\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|A_{i}z|^{2}\,dt\leq\mathbb{E}\int_{Q}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt.$
Then, combining (B.13)-(B.15), we obtain the following inequality for first term of the $J(w)$:
(B.16) $\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}&|rD_{i}w|^{2}\,dt\leq\;\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\\ &\mathbb{E}\int_{Q}s^{3}\varphi^{3}\lambda^{4}|z|^{2}\,dt+\mathbb{E}\int_{Q}s\mathcal{O}_{\lambda}((sh)^{2})|z|^{2}\,dt+\mathbb{E}\int_{Q}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt.\end{split}$
@@
Applying (1.15) to the first term on the right side of the previous equation, we obtain the following equality
(B.17) $rA_{i}D_{i}w=rA_{i}D_{i}(\rho)\,z+rA_{i}^{2}(\rho)\,A_{i}D_{i}(z)+\frac{h^{2}}{4}\left(rD_{i}^{2}(\rho)\,D_{i}A_{i}(z)+2rD_{i}A_{i}(\rho)\,D_{i}^{2}(z)\right).$
With Proposition A.3 and Lemma A.1 we then find
(B.18) $rA_{i}D_{i}w=A_{i}D_{i}z-s\lambda\varphi\mathcal{O}(1)\,z+\mathcal{O}_{\lambda}((sh)^{2})\,D_{i}A_{i}z+h\mathcal{O}_{\lambda}(sh)\,D_{i}^{2}(z)$
Now, by applying the triangle inequality and (B.18), we obtain the following estimate for each $i\in\llbracket 1,d\rrbracket$ of the second terms from $J(w)$:
(B.19) $\begin{split}\mathbb{E}\int_{Q}s\lambda^{2}\varphi&|rA_{i}D_{i}w|^{2}\,dt\\ \leq&\,\mathbb{E}\int_{Q}s\lambda^{2}\varphi|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}s^{3}\lambda^{4}\varphi\mathcal{O}(1)\,|z|^{2}\,dt+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}((sh)^{4})\,|D^{2}_{i}z|^{2}\,dt.\end{split}$
Combining (B.16) and (B.19), we can estimate to $J(w)$ as
(B.20) $\begin{split}J(w)\leq\sum_{i=1}^{n}&\left[\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{2}\varphi|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\right.\\ &\left.+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}((sh)^{4})\,|D^{2}_{i}z|^{2}\,dt\right]+\mathbb{E}\int_{Q}s^{3}\lambda^{4}\varphi\mathcal{O}(1)\,|z|^{2}\,dt.\end{split}$
@@
Now, by (A.4), (A.1) and recall that $r\rho=1$ it follows that
@@
Therefore, from the above equation for each $i\in\llbracket 1,n\rrbracket$, we obtain that
@@
Similar to (B.14) and (B.15), we find that
(B.21) $\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}\leq&\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|rD_{i}(w)|^{2}+\mathbb{E}\int_{Q}s^{2}\lambda^{2}\varphi^{2}(\partial_{i}\psi)^{2}|dz|^{2}\\ &+\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{4})|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}(sh)^{4})|dz|^{2}.\end{split}$
@@
Then, from (B.21) we obtain for each $i\in\llbracket 1,n\rrbracket$
(B.22) $\begin{split}&\mathbb{E}\int_{Q}s\lambda^{2}\varphi\gamma_{i}(\partial_{i}\psi)^{2}|rdw|^{2}-\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|rD_{i}(w)|^{2}\leq-\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}\\ &+\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{4})|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}\left(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi\right)\gamma_{i}(\partial_{i}\psi)^{2}|dz|^{2}+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}(sh)^{4})|dz|^{2}.\end{split}$
Du to the inequalities (B.20),(B.28) and (B.22), we conclude the proof of Lemma (B.7).This shows that for $0<sh\leq\varepsilon$ and $\lambda=\lambda_{1}$, there exist $\varepsilon_{2}(\lambda)>0$ and $s_{2}\geq 1$ such that for $s\geq s_{2}$, $0<h<h_{2},$ and $0<sh\leq\varepsilon_{2}(\lambda)$, the desired inequality is obtained. ∎
To proceed, we introduce the following lemma, which allows us to incorporate the second-order spatial term $D_{ij}^{2}$ into the left-hand side of the Carleman estimate. The proof follows by adapting the approach in (lecaros2025lipschitzstabilityinverseproblems, , Theorem 1.2).
Lemma B.8.
There exists $\varepsilon_{3}(\lambda)>0$ and $s_{3}\geq 1$ such that for $s\geq s_{3}$, $0<h<h_{3}$, and $0<sh\leq\varepsilon_{3}(\lambda)$, we obtain the following inequality:
$\begin{split}&\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}}s\lambda^{2}\varphi^{2}|\nabla\psi|_{\gamma}^{2}\gamma_{i}\partial_{i}\psi\,|rD_{i}w|^{2}+\mathcal{O}(s^{-1})|rA_{i}D_{i}w|^{2}\,dt\\ &\leq C_{\varepsilon_{3},s_{3}}\left(\mathbb{E}\int_{Q}|Cz|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{2}|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q}s\lambda^{2}\varphi^{2}|\nabla_{\gamma}\psi|^{2}\,|A_{i}D_{i}z|^{2}\,dt\right)\end{split}$
@@
Recalling that $r\sum_{i\in\llbracket 1,n\rrbracket}D_{i}(\gamma_{i}D_{i}w)=Cz+B_{2}z+B_{3}z-M_{h}z\triangleq\mathcal{B}w$ and applying Young’s inequality, we obtain
(B.23) $|\mathcal{B}w|^{2}\leq 2(|Cz|^{2}+|B_{2}z|^{2}+|B_{3}z|^{2}+|M_{h}z|^{2})$
Therefore, taking $s>1$ we have a first estimate for the third left-side term of (B.12) given by
(B.24) $\begin{split}\mathbb{E}\int_{Q}s^{-1}|\mathcal{B}w|^{2}\,dt\leq\mathbb{E}\int_{Q}2|Cz|^{2}\,dt+&\mathbb{E}\int_{Q}2s^{-1}|B_{2}z|^{2}\,dt\\ &+\mathbb{E}\int_{Q}2s^{-1}|B_{3}z|^{2}\,dt+\mathbb{E}\int_{Q}2s^{-1}|M_{h}z|^{2}\,dt.\end{split}$
From the definition of the $B_{2}z$ and $B_{3}z$ respectively, Lemma A.1 and Proposition A.3, we can see that
(B.25) $\begin{split}\mathbb{E}\int_{Q}2s^{-1}|B_{2}z|^{2}dt\leq&\sum_{i=1}^{n}\mathbb{E}\int_{Q}8s^{-1}|\gamma_{i}rD_{i}A_{i}\rho\,D_{i}A_{i}z|^{2}dt\\ \leq&8\sum_{i=1}^{n}\mathbb{E}\int_{Q}(s\lambda^{2}\varphi^{2}|\nabla_{\gamma}\psi|^{2}+\mathcal{O}_{\lambda}|((sh)^{4})||A_{i}D_{i}z|^{2}dt,\end{split}$
(B.26) $\mathbb{E}\int_{Q}2s^{-1}|B_{3}z|^{2}dt=4\mathbb{E}\int_{Q}2s^{-1}|s\Delta_{\gamma_{i}}\varphi\,z|^{2}dt\leq 8\mathbb{E}\int_{Q}(s\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{4}+s\lambda|\mathcal{O}((1)|^{2}|z|^{2}\;dt$
and similar to Lemma B.5, we can deduce the following estimate for $M_{h}z$ as
(B.27) $\mathbb{E}\int_{Q}2s^{-1}|M_{h}z|^{2}\,dt\leq\mathcal{O}_{\lambda}(1)\left(\mathbb{E}\int_{Q}s|z|^{2}\,dt+h^{2}\sum_{i=1}^{n}\int_{Q^{\ast}}s|D_{i}z|^{2}\,dt\right).$
Thus, combining (B.24)-(B.27) we obtain the estimate
(B.28) $\begin{split}\mathbb{E}\int_{Q}&s^{-1}|\mathcal{B}w|^{2}\,dt\\ \leq&\;\mathbb{E}\int_{Q}|Cz|^{2}dt+\mathbb{E}\int_{Q}(4s\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{4}+s\mathcal{O}_{\lambda}(1)+s\lambda|\mathcal{O}((1)|^{2}|z|^{2}dt\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q}h\mathcal{O}_{\lambda}(sh)|D_{i}z|^{2}dt+\mathbb{E}\int_{Q}(4s\lambda^{2}\varphi^{2}|\nabla_{\gamma}\psi|^{2}+\mathcal{O}_{\lambda}|((sh)^{4})||A_{i}D_{i}z|^{2}dt\right).\end{split}$
@@
$\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{*}}s^{-1}\gamma_{i}\gamma_{j}r^{2}|D_{ij}w|^{2}\;dt$. To this end, we can now proceed analogously to the proof of Theorem 1.2 in lecaros2025lipschitzstabilityinverseproblems . In particular, from the inequalities (2.27)-(2.32) in lecaros2025lipschitzstabilityinverseproblems it follows that
(B.29) $\begin{split}\mathbb{E}\int_{Q}s^{-1}|\mathcal{B}w|^{2}\,dt\geq&\sum_{i,j\in\llbracket 1,n\rrbracket}\int_{Q_{ij}^{\ast}}s^{-1}\gamma_{i}\gamma_{j}|rD_{ij}^{2}w|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}}s\lambda^{2}\varphi^{2}|\nabla\psi|_{\gamma}^{2}\gamma_{i}\partial_{i}\psi|rD_{i}w|^{2}\,dt\\ &-\sum_{i=1}^{n}\mathbb{E}\int_{Q}s^{-1}r^{2}|\mathcal{O}(1)|^{2}|A_{i}D_{i}w|^{2}\,dt-K(w),\end{split}$

### v2

@@ A.3. Returning to the original variable
(A.12) $\begin{split}J(w)+&\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}|rw|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{2}_{1}\varphi|rg|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}|rD_{i}(g)|^{2}\,dt\\ &\leq\mathcal{C}^{\prime}_{\tau_{2},\varepsilon_{2}}\left(\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}|z|^{2}\,dt+\mathbb{E}\int_{Q}(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi)|\nabla_{\gamma}\psi|^{2}|dz|^{2}\right.\\ &\left.+\sum_{i=1}^{n}\left[\mathbb{E}\int_{Q}s\lambda^{2}\varphi|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}|D_{i}(dz)|^{2}\right]\right)\end{split}$
where $J(w):=\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}|rD_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{2}\varphi|rA_{i}D_{i}w|^{2}\,dt$ and $w$ is solution of $dw+\sum_{i=1}^{n}D_{i}(\gamma_{i}w)dt=fdt+gdB(t)$ with $w=0$ on $\partial Q$.
@@
Recalling $w=\rho z$, we have $D_{i}w=A_{i}\rho\,D_{i}z+D_{i}\rho\,A_{i}z$. Therefore, by (Boyer et al., 2010b, Lemma 3.7) and (Pérez, 2024, Proposition 3.3), for each i=1,,n.i{=1,...,n.} it holds that
(A.13) $\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}|rD_{i}w|^{2}\,dt\leq\;&\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\\ &\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\varphi^{3}\lambda^{4}|A_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|A_{i}z|^{2}\,dt.\end{split}$
@@
(A.14) $\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\varphi^{3}\lambda^{4}|A_{i}z|^{2}\,dt\leq\mathbb{E}\int_{Q}s^{3}\varphi^{3}\lambda^{4}|z|^{2}\,dt+\mathbb{E}\int_{Q}s\mathcal{O}_{\lambda}((sh)^{2})|z|^{2}\,dt.$
@@
(A.15) $\mathbb{E}\int_{Q_{i}^{\ast}}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|A_{i}z|^{2}\,dt\leq\mathbb{E}\int_{Q}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt.$
Then, combining (A.13)-(A.15), we obtain the following inequality for first term of the $J(w)$:
(A.16) $\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}&|rD_{i}w|^{2}\,dt\leq\;\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\\ &\mathbb{E}\int_{Q}s^{3}\varphi^{3}\lambda^{4}|z|^{2}\,dt+\mathbb{E}\int_{Q}s\mathcal{O}_{\lambda}((sh)^{2})|z|^{2}\,dt+\mathbb{E}\int_{Q}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt.\end{split}$
@@
Applying (1.16) to the first term on the right-hand side of the previous equation, we obtain the following equality
(A.17) $rA_{i}D_{i}w=rA_{i}D_{i}(\rho)\,z+rA_{i}^{2}(\rho)\,A_{i}D_{i}(z)+\frac{h^{2}}{4}\left(rD_{i}^{2}(\rho)\,D_{i}A_{i}(z)+2rD_{i}A_{i}(\rho)\,D_{i}^{2}(z)\right).$
With Proposition (Pérez, 2024, Proposition 3.3) and (Boyer et al., 2010b, Lemma 3.7) we then find
(A.18) $rA_{i}D_{i}w=A_{i}D_{i}z-s\lambda\varphi\mathcal{O}(1)\,z+\mathcal{O}_{\lambda}((sh)^{2})\,D_{i}A_{i}z+h\mathcal{O}_{\lambda}(sh)\,D_{i}^{2}(z)$
Now, by applying the triangle inequality and (A.18), we obtain the following estimate for each i=1,,n.i{=1,...,n.} of the second terms from $J(w)$:
(A.19) $\begin{split}\mathbb{E}\int_{Q}s\lambda^{2}\varphi&|rA_{i}D_{i}w|^{2}\,dt\\ \leq&\,\mathbb{E}\int_{Q}s\lambda^{2}\varphi|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}s^{3}\lambda^{4}\varphi\mathcal{O}(1)\,|z|^{2}\,dt+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}((sh)^{4})\,|D^{2}_{i}z|^{2}\,dt.\end{split}$
Combining (A.16) and (A.19), we can estimate to $J(w)$ as
(A.20) $\begin{split}J(w)\leq\sum_{i=1}^{n}&\left[\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi\lambda^{2}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{2}\varphi|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\right.\\ &\left.+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}((sh)^{4})\,|D^{2}_{i}z|^{2}\,dt\right]+\mathbb{E}\int_{Q}s^{3}\lambda^{4}\varphi\mathcal{O}(1)\,|z|^{2}\,dt.\end{split}$
@@
Now, by (Boyer et al., 2010b, Lemma 3.7), (Boyer et al., 2010b, Corollary 3.8) and recall that $r\rho=1$ it follows that
@@
Therefore, from the above equation for each i=1,,n.i{=1,...,n.}, we obtain that
@@
Similar to (A.14) and (A.15), we find that
(A.21) $\begin{split}\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}\leq&\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|rD_{i}(w)|^{2}+\mathbb{E}\int_{Q}s^{2}\lambda^{2}\varphi^{2}(\partial_{i}\psi)^{2}|dz|^{2}\\ &+\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{4})|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}(sh)^{4})|dz|^{2}.\end{split}$
@@
Then, from (A.21) we obtain for each i=1,,n.i{=1,...,n.}
(A.22) $\begin{split}&\mathbb{E}\int_{Q}s\lambda^{2}\varphi\gamma_{i}(\partial_{i}\psi)^{2}|rdw|^{2}-\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|rD_{i}(w)|^{2}\leq-\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}\\ &+\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{4})|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}\left(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi\right)\gamma_{i}(\partial_{i}\psi)^{2}|dz|^{2}+\mathbb{E}\int_{Q}\mathcal{O}_{\lambda}(sh)^{4})|dz|^{2}.\end{split}$
Du to the inequalities (A.20),(A.28) and (A.22), we conclude the proof of Lemma (A.7).This shows that for $0<sh\leq\varepsilon$ and $\lambda=\lambda_{1}$, there exist $\varepsilon_{2}(\lambda)>0$ and $s_{2}\geq 1$ such that for $s\geq s_{2}$, $0<h<h_{2},$ and $0<sh\leq\varepsilon_{2}(\lambda)$, the desired inequality is obtained. ∎
To proceed, we introduce the following lemma, which allows us to incorporate the second-order spatial term $D_{ij}^{2}$ into the left-hand side of the Carleman estimate. The proof follows by adapting the approach in (Lecaros et al., 2025, Theorem 1.2).
Lemma A.8.
There exists $\varepsilon_{3}(\lambda)>0$ and $\tau_{3}\geq 1$ such that for $s\geq s_{3}$, $0<h<h_{3}$, and $0<sh\leq\varepsilon_{3}(\lambda)$, we obtain the following inequality:
$\begin{split}&\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}}s\lambda^{2}\varphi^{2}|\nabla\psi|_{\gamma}^{2}\gamma_{i}\partial_{i}\psi\,|rD_{i}w|^{2}+\mathcal{O}(s^{-1})|rA_{i}D_{i}w|^{2}\,dt\\ &\leq\mathcal{C}_{\varepsilon_{3},\tau_{3}}\left(\mathbb{E}\int_{Q}|Cz|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{2}|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q}s\lambda^{2}\varphi^{2}|\nabla_{\gamma}\psi|^{2}\,|A_{i}D_{i}z|^{2}\,dt\right)\end{split}$
@@
Recalling that $r{\sum_{i=1}^{n}}D_{i}(\gamma_{i}D_{i}w)=Cz+B_{2}z+B_{3}z-M_{h}z:=\mathcal{B}w$ and applying Young’s inequality, we obtain
(A.23) $|\mathcal{B}w|^{2}\leq 2(|Cz|^{2}+|B_{2}z|^{2}+|B_{3}z|^{2}+|M_{h}z|^{2})$
Therefore, taking $s>1$ we have a first estimate for the third left-side term of (A.12) given by
(A.24) $\begin{split}\mathbb{E}\int_{Q}s^{-1}|\mathcal{B}w|^{2}\,dt\leq\mathbb{E}\int_{Q}2|Cz|^{2}\,dt+&\mathbb{E}\int_{Q}2s^{-1}|B_{2}z|^{2}\,dt\\ &+\mathbb{E}\int_{Q}2s^{-1}|B_{3}z|^{2}\,dt+\mathbb{E}\int_{Q}2s^{-1}|M_{h}z|^{2}\,dt.\end{split}$
From the definition of the $B_{2}z$ and $B_{3}z$ respectively, (Boyer et al., 2010b, Lemma 3.7) and Proposition (Pérez, 2024, Proposition 3.3), we can see that
(A.25) $\begin{split}\mathbb{E}\int_{Q}2s^{-1}|B_{2}z|^{2}dt\leq&\sum_{i=1}^{n}\mathbb{E}\int_{Q}8s^{-1}|\gamma_{i}rD_{i}A_{i}\rho\,D_{i}A_{i}z|^{2}dt\\ \leq&8\sum_{i=1}^{n}\mathbb{E}\int_{Q}(s\lambda^{2}\varphi^{2}|\nabla_{\gamma}\psi|^{2}+\mathcal{O}_{\lambda}|((sh)^{4})||A_{i}D_{i}z|^{2}dt,\end{split}$
(A.26) $\mathbb{E}\int_{Q}2s^{-1}|B_{3}z|^{2}dt=4\mathbb{E}\int_{Q}2s^{-1}|s\Delta_{\gamma_{i}}\varphi\,z|^{2}dt\leq 8\mathbb{E}\int_{Q}(s\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{4}+s\lambda|\mathcal{O}((1)|^{2}|z|^{2}\;dt$
and similar to Lemma A.5, we can deduce the following estimate for $M_{h}z$ as
(A.27) $\mathbb{E}\int_{Q}2s^{-1}|M_{h}z|^{2}\,dt\leq\mathcal{O}_{\lambda}(1)\left(\mathbb{E}\int_{Q}s|z|^{2}\,dt+h^{2}\sum_{i=1}^{n}\int_{Q^{\ast}}s|D_{i}z|^{2}\,dt\right).$
Thus, combining (A.24)-(A.27) we obtain the estimate
(A.28) $\begin{split}\mathbb{E}\int_{Q}&s^{-1}|\mathcal{B}w|^{2}\,dt\\ \leq&\;\mathbb{E}\int_{Q}|Cz|^{2}dt+\mathbb{E}\int_{Q}(4s\lambda^{4}\varphi^{2}|\nabla_{\gamma}\psi|^{4}+s\mathcal{O}_{\lambda}(1)+s\lambda|\mathcal{O}((1)|^{2}|z|^{2}dt\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q}h\mathcal{O}_{\lambda}(sh)|D_{i}z|^{2}dt+\mathbb{E}\int_{Q}(4s\lambda^{2}\varphi^{2}|\nabla_{\gamma}\psi|^{2}+\mathcal{O}_{\lambda}|((sh)^{4})||A_{i}D_{i}z|^{2}dt\right).\end{split}$
@@
$\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{*}}s^{-1}\gamma_{i}\gamma_{j}r^{2}|D_{ij}w|^{2}\;dt$. To this end, we can now proceed analogously to the proof of Theorem 1.2 in Lecaros et al. (2025). In particular, from the inequalities (2.27)-(2.32) in Lecaros et al. (2025) it follows that
(A.29) $\begin{split}\mathbb{E}\int_{Q}s^{-1}|\mathcal{B}w|^{2}\,dt\geq&{\sum_{i,j=1}^{n}}\int_{Q_{ij}^{\ast}}s^{-1}\gamma_{i}\gamma_{j}|rD_{ij}^{2}w|^{2}\,dt-\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}}s\lambda^{2}\varphi^{2}|\nabla\psi|_{\gamma}^{2}\gamma_{i}\partial_{i}\psi|rD_{i}w|^{2}\,dt\\ &-\sum_{i=1}^{n}\mathbb{E}\int_{Q}s^{-1}r^{2}|\mathcal{O}(1)|^{2}|A_{i}D_{i}w|^{2}\,dt-K(w),\end{split}$